1. What is the reducing agent in the reaction below? Cr2O+ 3NO2+8H+-------> 2Cr3+ +3NO+ 4H3O

2. based on the reaction below, how many grams of vitamin-C (C6H2O4)) is present in a sample if 14.55ml of 0.1 M KIO3 is used during the titration?

IO2(aq)+ 8I (aq)+6H^+-------->3I3(aq) + 3H2O(l)

C6H7O2(aq)+ I3(aq) + H2O(l)--------> C4H7O(aq) + 3I (aq) +2H+(aq)

a.0.256g

b.0.085g

c.0.0146g

d. 0.769g
Answer #1

1) The reducing agent is the one that oxidizes, it is NO2.

2) The moles of citric acid present are calculated:

n CA = M * V * 3 = 0.1 M * 0.01455 L * 3 = 0.004 mol

The mass is calculated:

m = n * MM = 0.004365 * 192.12 = 0.769 g

Option D.


If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.